If X follows a normal distribution with u = 30 and o=4, what is P(X<27)?


1
Expert's answer
2022-03-27T11:00:00-0400

"P\left( X<27 \right) =P\left( \frac{X-30}{4}<\frac{27-30}{4} \right) =P\left( Z<-0.75 \right) =\varPhi \left( -0.75 \right) =0.2266"